OpenStudy (anonymous):

Which of the following is the solution of log (x + 5)0.001 = -3. (x+5)^-3=0.001. where would i go from here?

OpenStudy (anonymous):

note that \[0.001=10^{-3}\]

OpenStudy (anonymous):

making \(x+5=10\)

OpenStudy (anonymous):

how did you get to x+5=10

OpenStudy (anonymous):

i could rewrite it as (x+5)^-3=10^-2? how do i get rid of the exponents though

OpenStudy (anonymous):

\[(x+5)^{-3}=10^{-3}\iff x+5=10\]

OpenStudy (anonymous):

but one has an exponent of -3 and the other has -2? do do htye cancel?

OpenStudy (anonymous):

\(0.001=10^{-3}\) not \(10^{-2}\) \[10^{-2}=0.01\]

OpenStudy (anonymous):

oh i see thank you. that makes sense